OpenStudy (anonymous):

1/42 over 1/6 = 1/6 over x, x equals

OpenStudy (anonymous):

\[\frac{1}{46}\times \frac{6}{1}=\frac{1}{6x}\]

OpenStudy (anonymous):

oh wait \[\frac{\frac{1}{6}}{x}\]?

OpenStudy (anonymous):

is that the right hand side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it is \[\frac{6}{46}=\frac{\frac{1}{6}}{x}\]

OpenStudy (anonymous):

then solve as a ratio as \[6x=\frac{1}{6}\times 46\] and then divide by 6 to get \[x=\frac{46}{36}=\frac{23}{18}\]

OpenStudy (anonymous):

still a little lost but thanks for your help

OpenStudy (anonymous):

we can go slower if you like

OpenStudy (anonymous):

lets start with \[\frac{\frac{1}{46}}{\frac{1}{6}}\]

OpenStudy (anonymous):

oh damn! i have been writing 46 but it is 42. now wonder you are confused!

OpenStudy (anonymous):

so answer is wrong but method is right. lets start again

OpenStudy (anonymous):

\[\frac{\frac{1}{42}}{\frac{1}{6}}=\frac{1}{42}\times \frac{6}{1}\] because that is how you divide fractions, by multiplying by the reciprocal

OpenStudy (anonymous):

so the left hand side of this equation is \[\frac{6}{42}=\frac{1}{7}\] reduced

OpenStudy (anonymous):

so we have now \[\frac{1}{7}=\frac{\frac{1}{6}}{x}\]

OpenStudy (anonymous):

and now we solve this as a ratio by "cross multiplying" so you actually get the answer is one step

OpenStudy (anonymous):

\[1\times x = \frac{1}{6}\times 7\]

OpenStudy (anonymous):

better written as \[x=\frac{7}{6}\] sorry about writing 46 instead of 42

OpenStudy (anonymous):

i see thanks a lot!!

OpenStudy (anonymous):

yw